# Lectures on Gauge theories and Many-Body systems

### v1

@@ 6.1 String theory perspective
Noncommutative gauge theories [Douglas:2001ba] require string theory for their definition. In string theory realization of gauge theory, typically, there is an infinite number of degrees of freedom involved, of which only a finite number is relevant for low energy physics. Specifically, the gauge origami [Nekrasov:2016ydq] setup embeds the four dimensional spacetime ${\mathbb{C}}^{2}\approx{\mathbb{R}}^{4}$ as a subspace of ${\mathbb{C}}^{4}\approx{\mathbb{R}}^{8}$. The ten dimensional space-time of Type IIB string theory used in many of these constructions, is ${\mathbb{R}}^{8}\times{\mathbb{T}}^{2}$. Metrically, it is the twisted product, i.e. the total space of rank $4$ vector bundle over ${\mathbb{T}}^{2}$ with the $SU(4)$ structure group. As a complex manifold it is the total space of direct sum of four line bundles $L_{1},L_{2},L_{3},L_{4}$ over an elliptic curve $E$, such that the tensor product $L_{1}\otimes L_{2}\otimes L_{3}\otimes L_{4}\approx{\mathcal{O}}_{E}$. The parameters $L_{1},L_{2},L_{3},L_{4}\in Pic_{0}(E)$ of this background ${{\varepsilon}}_{1},{{\varepsilon}}_{2},{{\varepsilon}}_{3},{{\varepsilon}}_{4}$ are usually viewed additively, so
@@
Of course, one can get a more general flat ten dimensional metrics, by using a generic flat $SO(8)$ connection on $T^{2}$. The latter abelianizes to the maximal torus $U(1)^{4}$ of $SO(8)$. The significance of the reduction to $SU(4)$ with the maximal torus $U(1)^{3}\subset U(1)^{4}$ is the preservation of some nilpotent symmetry, supersymmetry. Mathematically this is saying that we have an infinite dimensional complex of (string) fields, whose cohomology is of interest to us.
@@
Macroscopically, the theory on the stack of $D5$ branes is four dimensional, and at low energy it should be described by a deformed version of ${\mathcal{N}}=4$ super-Yang-Mills theory with gauge group $U(N)$. The complexified coupling of this gauge theory is
@@
One can study various limits where the torus $T^{2}$ shrinks. It could shrink to $S^{1}$ or to a point. One scales the dilaton field $\phi$ to keep the effective gauge coupling in five or four dimensions finite. Mathematically, the theory on finite $T^{2}$ corresponds to elliptic cohomology of the moduli space of instantons on ${\mathbf{R}}^{4}$, the theory on $T^{2}$ collapsed to $S^{1}$ corresponds to $K$-theory, and the theory on $T^{2}$ collapsed to a point is the ordinary cohomology theory. Some details of this construction can be found in [Nekrasov:1996cz, Baulieu:1997nj].
@@ 6.2 Partitions
Partition is a set of non increasing integer numbers:
@@
the outer bounary of $\lambda$ is
@@ 6.3 Characters
Another common notation for such monomials is $q_{1}^{i-1}q_{2}^{j-1}=e^{c(\square=(i,j))}$.
@@
Let $M$ is a $\mathbb{C}^{*}\times\mathbb{C}^{*}$-module. It is just a $\mathbb{C}$-vector space equipped with a linear action of $\mathbb{C}^{*}\times\mathbb{C}^{*}\ni(q_{1},q_{2})$. It has a decomposition $M=\bigoplus\limits_{(i,j)\in\mathbb{Z}^{2}}M^{(i,j)}$, where $M^{(i,j)}=\{m\in M:(q_{1},q_{2}).m=q_{1}^{i-1}q_{2}^{j-1}m\}$. Let $M^{(i,j)}=0$ for $i<<0$ or $j<<0$. The character of $M$ is a formal power series
$\mathrm{ch}M=\sum\limits_{(i,j)\in\mathbb{Z}^{2}}\dim M^{(i,j)}q_{1}^{i-1}q_{2}^{j-1}\in\mathbb{C}((q_{1},q_{2})).$
@@
$\mathrm{ch}({M}_{\bullet})=\sum\limits_{k\in\mathbb{Z}}(-1)^{k}\mathrm{ch}({M}_{k})\in\mathbb{C}((q_{1},q_{2})).$ (6.27)
@@
Let $d_{k}:M_{k}\rightarrow M_{k+1}$ be the differentials of $M_{\bullet}$. We have an exact sequences
@@ 6.4 Measures
Four dimensional gauge theories have special solutions of equations of motions (critical points of the action) called instantons ([Nekrasov:2005wp], [FRENKEL2007215]). As we explained in the introduction to this section, equivariant localization of path integrals ([Nekrasov:2002qd]) done in two steps: first to instanton moduli spaces, then to fixed points of global symmetry acting on the moduli space produces interesting measures $\mu=\mu_{(\lambda^{(\alpha)})}$ on the set of multipartitions
@@ 6.5 Examples based on Dynkin diagrams
For example, the ${\hat{A}}_{r}$-theory has $W={\emptyset}$,
@@
Let ${\bf N}=\{1,\ldots,N\}$. For reasons explained in [Nekrasov2015BPSCFTCN] the set $A$ is the direct product $A=V\times{\bf N}$, with the map $\nu$} being the projection onto the first factor.
@@
Here, the set $E$ contains $_{i}{\rightarrow}_{j}$ where either $j=i+1$ for $i=1,\ldots,r-2$, or $i=0$ and $j=2$, or i=r2i=r-2 and $j=r$. Let ${\bf N}_{2}=\{1,\ldots,2N\}\approx{\bf N}\amalg{\bf N}$.For reasons explained in [Nekrasov2015BPSCFTCN] the set $A$ is the union of the direct products $A=\{0,1,r-1,r\}\times{\bf N}\amalg\{2,\ldots,r-2\}\times{\bf N}_{2}$, with the map $\nu$} being the projection onto the first factor in each summand.
@@ 6.6 Three theories
In agreement with the theory [Arnold:1997] of mathematical trinities there are three versions of everything. In our case we have the cohomological $H$, $K$-theoretic $K$ and elliptic cohomology $Ell$ versions of the model. The gauge theoretic origin of these models are the four dimensional gauge theory, five dimensional theory compactified on a circle, or the six dimensional theory compactified on a two-torus.
@@
$\mathfrak{p}=e^{2\pi\sqrt{-1}\sigma}\ .$
Depending on the type of cohomology theory we use the measure is expressed with the help of the function $\theta(x)$ defined as:
@@
Definition 6.6.
6.7 Example: $\hat{A}_{0}$-theory
Let us present in detail the example of $\hat{A}_{0}$-theory. It is the quiver model, with the both sets $V=\{{\bullet}\}$ and $E=\{{\rightarrow}\}$ consisting of one element, with the maps $s({\rightarrow})=t({\rightarrow})={\bullet}$. The set $A=\{1,\ldots,N\}$.
The index set $A$ in that case is just $\{1,...,N\}$. Its elements are called colors. The parameters of the measure depends are
@@ 6.7 Example: $\hat{A}_{0}$-theory
$q_{i}=e^{\varepsilon_{i}}$
$\mathfrak{q}_{\bullet}={\mathfrak{q}}=e^{2\pi\sqrt{-1}\tau},\penalty 10000\ {\operatorname{Im}\,}\tau>0.$
@@
where $\int\limits_{\mathrm{Hilb}^{[n]}(\mathbb{C}^{2})}$ is a pushforward to the point in appropriate equivariant cohomology theory ([OkoPan2004]).
@@
In the 4d case with $\theta(x)=x$ this formula reduces [Nekrasov2015BPSCFTCN] to
@@
$N=2$ partition function $Z^{N=2}$ for cohomological case describes a conformal block for Liouville CFT (++ free boson) ([Alday:2009aq]) on pointed elliptic curve $C_{\tau}$.
@@ 6.8 Vershik-Kerov limit
Now take the limit $\hbar\rightarrow 0$ with finite $\Lambda$. In this limit one finds the study of the limit shape of very large Young diagrams due to A. M. Vershik and S. V. Kerov ([Logan1977AVP], [VerKer77], [VerKer85]).
@@ 7 Order and Disorder
The view on quantum field theory as a limit of statistical mechanics suggests two types of interesting observables. The first type, the so-called order operators, are defined in terms of elementary fields locally, i.e. they are renormalized finite expressions in fields and their derivatives, evaluated at a point of spacetime, on a loop, a surface etc., the main feature being the locality of the associated functional. The next class of observables, the disorder operators, are defined in a tricky way. Sometimes they can also be expressed through the elementary fields of the theory, but in a non-local way. Typically one defines them as a prescription to perform the path integration over the space of fields, which are well-defined outside a submanifold $N\subset M$ of spacetime manifold, and have a singularity along $N$ of prescribed type, or asymptotics. In the following we consider examples of these operators appearing in particular quantum field theories. Then we give a hint on how equivariant localization discussed in (6) turns order and disorder observables in four-dimensional supersymmetric gauge theories to specific functions on the set of multipartitions. We don’t give a coherent derivation though but rather introduce the final answer for these functions as a definition. Then we study their expectation values.
@@ 7.1 Abelian sigma model
where ${\varphi}(x)\in V$ is a lift of the value of the field at $x$ to $V$ from $T$, and :::: is the sign of the normal ordering. In quantum theory the field $\varphi(x)$ does not really have a value at the point, however its circle average
@@
over the small circle of radius $\epsilon$} centered at $x$, has. So,
@@ 7.2 Wilson loop in gauge theory
where ${\Phi}(t)=\sum_{I=1}^{n}{\phi}^{I}{\dot{y}}_{I}(t)dt$, where $I$ labels the adjoint scalars in the theory (the ${\mathcal{N}}=4$ super-Yang-Mills theory in four dimensions has $n=6$, ${\mathcal{N}}=2$ has $n=2$) and ${\vec{y}}(t)$ is a parametrized loop in the auxiliary space ${\mathbb{R}}^{n}$. One can also include adjoint fermions via
@@ 7.5 Wilson loop in non-abelian gauge theory
The coadjoint orbits ${\mathcal{O}}_{\xi}$ are the symplectic leaves of Kirillov-Konstant Poisson structure on $\mathfrak{g}$$^{*}$:
@@
One way to understand (7.29) is to use an infinite-dimensional version of Duistermaat-Heckman formula (DH) [DH0, DH1, DH2]. Consider the space $L{\mathcal{O}}_{\xi}$ of parametrized loops ${\eta}:S^{1}\to{\mathcal{O}}_{\xi}$. It is an infinite dimensional symplectic manifold with the symplectic form given by the average pullback of ${\omega}_{\xi}$:
@@
We are therefore in the realm of the DH formula in the formulation of [DH2]: let $\mathbb{T}$} be the torus of $\widetilde{LG}$} generated by $\phi$. By conjugation (Floquet-Lyapounov theory) $\phi$} can be brought to the form
@@
Assuming (without loss of generality) that $\xi\in{\mathfrak{t}}^{*}$, we see that these intersection points are parametrized with an elements of the Weyl group $W$:
@@
Evaluating the fixed contributions, using the $\zeta$-function regularization of infinite products – the denominator of DH formula involves the product of weights of $\mathbb{T}$}-action on $T_{w\xi}\left(L{\mathcal{O}}_{\xi}\right)$ – one arrives at Weyl character formula [HW].
Summarizing, the Wilson loop observable $W_{R}(C)$ in any gauge theory is a result of coupling of that gauge theory to an auxiliary quantum mechanical system, assigned to the loop $C$, which can be viewed as the worldline or periodic time of that quantum system.
7.6 Wilson loop in Chern-Simons theory from disorder operator
Chern-Simons theory is a three dimensional gauge theory with the gauge group $G$, the action $kS_{CS}$ for an integer $k$ and
@@ 7.6 Wilson loop in Chern-Simons theory from disorder operator
Now we insert $W_{R}(C)$ in Chern-Simons path integral:
@@
7.7 Codimension 2 vortex-type operators in four dimensional Yang-Mills theory
The example above of three dimensional Chern-Simons theory motivated introduction of operators like (7.43), but the really interesting playground for their study is the four dimensional (super)-Yang-Mills theory. One can define the operators/observables associated with two dimensional surfaces $\Sigma$ embedded into the four dimensional spacetime, as an instruction to perform the path integral over the space of gauge fields with a singularity along $\Sigma$, modelled on (7.43). Now, $\eta$ will vary along $\Sigma$, while remaining inside $\mathcal{O}_{\xi}$. In other words, we have coupled the four dimensional gauge theory to a two dimensional sigma model valued in ${\mathcal{O}}_{\xi}$. Unlike the case of loop operators in three dimensions, the two dimensional sigma model valued in ${\mathcal{O}}_{\xi}$ comes with a plethora of additional topological charges

### v2

@@ 6.1 String theory perspective
Noncommutative gauge theories [DN01] require string theory for their definition. In string theory realization of gauge theory, typically, there is an infinite number of degrees of freedom involved, of which only a finite number is relevant for low energy physics. Specifically, the gauge origami [Nek18] setup embeds the four dimensional spacetime ${\mathbb{C}}^{2}\approx{\mathbb{R}}^{4}$ as a subspace of ${\mathbb{C}}^{4}\approx{\mathbb{R}}^{8}$. The ten dimensional space-time of Type IIB string theory used in many of these constructions, is ${\mathbb{R}}^{8}\times{\mathbb{T}}^{2}$. Metrically, it is the twisted product, i.e. the total space of rank $4$ vector bundle over ${\mathbb{T}}^{2}$ with the $SU(4)$ structure group. As a complex manifold it is the total space of direct sum of four line bundles $L_{1},L_{2},L_{3},L_{4}$ over an elliptic curve $E$, such that the tensor product $L_{1}\otimes L_{2}\otimes L_{3}\otimes L_{4}\approx{\mathcal{O}}_{E}$. The parameters $L_{1},L_{2},L_{3},L_{4}\in Pic_{0}(E)$ of this background ${{\varepsilon}}_{1},{{\varepsilon}}_{2},{{\varepsilon}}_{3},{{\varepsilon}}_{4}$ are usually viewed additively, so
@@
Of course, one can get a more general flat ten dimensional metrics, by using a generic flat $SO(8)$ connection on $T^{2}$. The latter abelianizes to the maximal torus $U(1)^{4}$ of $SO(8)$. The significance of the reduction to $SU(4)$ with the maximal torus $U(1)^{3}\subset U(1)^{4}$ is the preservation of some nilpotent symmetry, supersymmetry. Mathematically this is saying that we have an infinite-dimensional complex of (string) fields, whose cohomology is of interest to us.
@@
Macroscopically, the theory on the stack of $D5$ branes is four dimensional, and at low energy it should be described by a deformed version of ${\mathcal{N}}=4$ super-Yang–Mills theory with gauge group $U(N)$. The complexified coupling of this gauge theory is
@@
One can study various limits where the torus $T^{2}$ shrinks. It could shrink to $S^{1}$ or to a point. One scales the dilaton field $\phi$ to keep the effective gauge coupling in five or four dimensions finite. Mathematically, the theory on finite $T^{2}$ corresponds to elliptic cohomology of the moduli space of instantons on ${\mathbf{R}}^{4}$, the theory on $T^{2}$ collapsed to $S^{1}$ corresponds to $K$-theory, and the theory on $T^{2}$ collapsed to a point is the ordinary cohomology theory. Some details of this construction can be found in [Nek98, BLN98].
@@ 6.2 Partitions
A partition is a set of non-increasing integer numbers:
@@
the outer boundary of $\lambda$ is
@@ 6.3 Characters
Another common notation for such monomials is $q_{1}^{i-1}q_{2}^{j-1}=e^{c(\square=(i,j))}$. Let ${\sf R}={\mathbb{C}}[[q_{1},q_{2}]][q_{1}^{-1},q_{2}^{-1}]$.
@@
Let $M$ be a $\mathbb{C}^{*}\times\mathbb{C}^{*}$-module. It is just a $\mathbb{C}$-vector space equipped with a linear action of $\mathbb{C}^{*}\times\mathbb{C}^{*}\ni(q_{1},q_{2})$. It has a decomposition $M=\bigoplus\limits_{(i,j)\in\mathbb{Z}^{2}}M^{(i,j)}$, where $M^{(i,j)}=\{m\in M:(q_{1},q_{2}).m=q_{1}^{i-1}q_{2}^{j-1}m\}$. Let $M^{(i,j)}=0$ for $i\ll 0$ or $j\ll 0$. The character of $M$ is a formal power series
$\mathrm{ch}M=\sum\limits_{(i,j)\in\mathbb{Z}^{2}}\dim M^{(i,j)}q_{1}^{i-1}q_{2}^{j-1}\in{\sf R}$
@@
$\mathrm{ch}({M}_{\bullet})=\sum\limits_{k\in\mathbb{Z}}(-1)^{k}\mathrm{ch}({M}_{k})\in{\sf R}.$ (6.27)
@@
Let $d_{k}:M_{k}\rightarrow M_{k+1}$ be the differentials of $M_{\bullet}$. We have exact sequences
@@ 6.4 Measures
Four dimensional gauge theories have special solutions of equations of motions (critical points of the action) called instantons ([Nek05], [FLN7]). As we explained in the introduction to this section, equivariant localization of path integrals ([Nek02]) done in two steps: first to instanton moduli spaces, then to fixed points of global symmetry acting on the moduli space produces interesting measures $\mu=\mu_{(\lambda^{(\alpha)})}$ on the set of multi–partitions
@@ 6.5 Examples based on Dynkin diagrams
For example, the $\widehat{{A}_{r}}-$theory has $W={\emptyset}$,
@@
Let ${\bf N}=\{1,\ldots,N\}$. For reasons explained in [Nek15] the set $A$ is the direct product $A=V\times{\bf N}$, with the map $\nu$} being the projection onto the first factor.
@@
Here, the set $E$ contains $_{i}{\rightarrow}_{j}$ where either $j=i+1$ for $i=1,\ldots,r-2$, or $i=0$ and $j=2$, or i=r2i=r-2 and $j=r$. Let ${\bf N}_{2}=\{1,\ldots,2N\}\approx{\bf N}\amalg{\bf N}$. For reasons explained in [Nek15] the set $A$ is the union of the direct products $A=\{0,1,r-1,r\}\times{\bf N}\amalg\{2,\ldots,r-2\}\times{\bf N}_{2}$, with the map $\nu$} being the projection onto the first factor in each summand.
@@ 6.6 Three theories
In agreement with the theory [Arn25] of mathematical trinities there are three versions of everything worth discussing. In our case we have the cohomological $H$, $K$-theoretic $K$, and elliptic cohomology $Ell$ versions of the model. These models originate in four-dimensional gauge theory, five-dimensional gauge theory compactified on a circle, and six-dimensional gauge theory compactified on a two-torus.
@@
$\mathfrak{p}=e^{2\pi{\mathrm{i}}\sigma}\ .$
Depending on the cohomology theory under consideration, the measure is expressed in terms of the function $\theta(x)$, defined by:
@@
6.7 Example: $\widehat{A_{0}}-$theory
Let us discuss the $\widehat{A_{0}}-$theory in detail. It is the quiver model in which both sets $V=\{{\bullet}\}$ and $E=\{{\rightarrow}\}$ consist of one element, with the maps $s({\rightarrow})=t({\rightarrow})={\bullet}$.
The index set $A$ in that case is just $\{1,...,N\}$. Its elements are called colors. The parameters of the measure are
@@ 6.7 Example: $\widehat{A_{0}}-$theory
$q_{a}=e^{\varepsilon_{a}}$
$\mathfrak{q}_{\bullet}={\mathfrak{q}}=e^{2\pi{\mathrm{i}}\tau},\penalty 10000\ {\rm Im}\tau>0.$
@@
where $\int\limits_{\mathrm{Hilb}^{[n]}(\mathbb{C}^{2})}$ is a pushforward to the point in appropriate equivariant cohomology theory ([OP04]).
@@
In the 4d case with $\theta(x)=x$ this formula reduces [Nek15] to
@@
$N=2$ partition function $Z^{N=2}$ for cohomological case describes a conformal block for Liouville CFT (++ free boson) ([AGT10]) on pointed elliptic curve $C_{\tau}$.
@@ 6.8 Vershik-Kerov limit
Now take the limit $\hbar\rightarrow 0$ with finite $\Lambda$. In this limit one finds the study of the limit shape of very large Young diagrams due to A. M. Vershik and S. V. Kerov ([LS77], [VK77], [VK85]).
@@ 7 Order and Disorder
The view on quantum field theory as a limit of statistical mechanics suggests two types of interesting observables. The first type, the so-called order operators, are defined in terms of elementary fields locally, i.e. they are renormalized finite expressions in fields and their derivatives, evaluated at a point of spacetime, on a loop, a surface etc., the main feature being the locality of the associated functional. The next class of observables, the disorder operators, are defined in a tricky way. Sometimes they can also be expressed through the elementary fields of the theory, but in a non-local way. Typically one defines them as a prescription to perform the path integration over the space of fields, which are well-defined outside a submanifold $N\subset M$ of spacetime manifold, and have a singularity along $N$ of prescribed type, or asymptotics. In the following we consider examples of these operators appearing in particular quantum field theories. Then we give a hint on how equivariant localization discussed in (6) turns order and disorder observables in four-dimensional supersymmetric gauge theories to specific functions on the set of multi–partitions. We do not derive these functions here; instead, we introduce their final expressions as definitions. Then we study their expectation values.
@@ 7.1 Abelian sigma model
where ${\varphi}(x)\in V$ is a lift of the value of the field at $x$ to $V$ from $T$, and :::: is the sign of the normal ordering. In quantum theory the field $\varphi(x)$ is not well-defined pointwise; its average
@@
over a small circle of radius $\epsilon$} centered at $x$, however, is well defined. So,
@@ 7.2 Wilson loop in gauge theory
where ${\Phi}(t)=\sum_{I=1}^{n}{\phi}^{I}{\dot{y}}_{I}(t)dt$, where $I$ labels the adjoint scalars in the theory (the ${\mathcal{N}}=4$ super-Yang–Mills theory in four dimensions has $n=6$, ${\mathcal{N}}=2$ has $n=2$) and ${\vec{y}}(t)$ is a parametrized loop in the auxiliary space ${\mathbb{R}}^{n}$. One can also include adjoint fermions via
@@ 7.5 Wilson loop in non-abelian gauge theory
The coadjoint orbits ${\mathcal{O}}_{\xi}$ are the symplectic leaves of the Kirillov–Kostant Poisson structure on $\mathfrak{g}$$^{*}$:
@@
One way to understand (7.29) is to use an infinite-dimensional version of Duistermaat-Heckman formula (DH) [Dui82, Ber82, Ati84]. Consider the space $L{\mathcal{O}}_{\xi}$ of parametrized loops ${\eta}:S^{1}\to{\mathcal{O}}_{\xi}$. It is an infinite-dimensional symplectic manifold with the symplectic form given by the average pullback of ${\omega}_{\xi}$:
@@
We are therefore in the realm of the DH formula in the formulation of [Ati84]: let $\mathbb{T}$} be the torus of $\widetilde{LG}$} generated by $\phi$. By conjugation (Floquet–Lyapunov theory) $\phi$} can be brought to the form
@@
Assuming (without loss of generality) that $\xi\in{\mathfrak{t}}^{*}$, we see that these intersection points are parametrized by elements of the Weyl group $W$:
@@
Evaluating the fixed contributions, using the $\zeta$-function regularization of infinite products – the denominator of DH formula involves the product of weights of $\mathbb{T}$}-action on $T_{w\xi}\left(L{\mathcal{O}}_{\xi}\right)$ – one arrives at Weyl character formula [Weyl].
In summary, the Wilson loop observable $W_{R}(C)$ in any gauge theory arises from coupling that theory to an auxiliary quantum–mechanical system assigned to the loop $C$.
7.6 Wilson loop in Chern–Simons theory from disorder operator
Chern–Simons theory is a three dimensional gauge theory with the gauge group $G$, the action $kS_{CS}$ for an integer $k$ and
@@ 7.6 Wilson loop in Chern–Simons theory from disorder operator
Now we insert $W_{R}(C)$ in Chern–Simons path integral:
@@
7.7 Codimension 2 vortex-type operators in four dimensional Yang–Mills theory
The three-dimensional Chern–Simons example above motivates the introduction of operators such as (7.43), but the most interesting setting for their study is the four dimensional (super)-Yang–Mills theory. One can define the operators/observables associated with two dimensional surfaces $\Sigma$ embedded into the four dimensional spacetime, as an instruction to perform the path integral over the space of gauge fields with a singularity along $\Sigma$, modelled on (7.43). Now, $\eta$ will vary along $\Sigma$, while remaining inside $\mathcal{O}_{\xi}$. In other words, we have coupled the four dimensional gauge theory to a two dimensional sigma model valued in ${\mathcal{O}}_{\xi}$. Unlike the case of loop operators in three dimensions, the two dimensional sigma model valued in ${\mathcal{O}}_{\xi}$ comes with a plethora of additional topological charges